OpenStudy (anonymous):

Simplify the expression. 2'4 A:16 B:8 C:8 D:16

OpenStudy (bibby):

\(2^4\)?

OpenStudy (anonymous):

B AND C

OpenStudy (anonymous):

yea @bibby

OpenStudy (anonymous):

If you're asking for \[2^4\] then do, 4*4*4*4 = 16

OpenStudy (bibby):

do you know how exponents work? \(5^3=5*5*5\)

OpenStudy (anonymous):

no but since you did that i think i got it

OpenStudy (bibby):

to confirm, it's 2*2*2*2 or 4*4 or 16

OpenStudy (anonymous):

wait how do i do this 12'0

OpenStudy (anonymous):

12

OpenStudy (bibby):

Most things to the 0 power are 1. I can't remember what 0^0 is

OpenStudy (bibby):

that means \(5^0=1\) \(11^0=1\) etc.

OpenStudy (anonymous):

thank you